A survey of 1273 student loan borrowers found that 458 had loans totaling more than $20,000 for their undergraduate education. Give a 99% confidence interval for the proportion of all student loan borrowers who have loans of $20,000 or more for their undergraduate education. (Round your answers to three decimal places.)

Lower limit _______
Upper limit _______
Answer #1

Solution :

Given that,

n = 1273

x = 458

Point estimate = sample proportion = \hat p = x / n = 458 / 1273 = 0.360

1 - \hat p = 1 - 0.360 = 0.64

At 99% confidence level the z is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 2.576 * (\sqrt((0.360 * 0.64) / 1273)

= 0.035

A 95% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.360 - 0.035< p < 0.360 + 0.035

0.325 < p < 0.395

( 0.325 , 0.395 )

Lower limit 0.325
Upper limit 0.395